OpenStudy (anonymous):

check answer... a club with 36 members..18 takes math 5 takes both math and literature and 8 take neither. how many take lit but not math.

OpenStudy (anonymous):

is it 5?

OpenStudy (anonymous):



OpenStudy (anonymous):

fill in the missing part with \(36-8-18\)

OpenStudy (anonymous):

lord i didnt think of doing it that way

OpenStudy (anonymous):

see what i did was just subtracted the numbers from 36. is that a right way or no?

OpenStudy (anonymous):

@satellite73 i think 36-8 = 28. :P @tena32 NOPE. Because when they say "Took math" , it includes people who took both math and lit as well. Hence people who took math but not lit = 18-5 = 13.

OpenStudy (anonymous):

ok so i do 18-5 first? im sorry if i dont understand this very well lol

OpenStudy (anonymous):

and that is 13...

OpenStudy (anonymous):

@tena32 ok. the thing is; there are 3 sets of people: 1. Who took ONLY MATH BUT NOT LIT. 2. Who took both math and lit. 3. Who took ONLY LIT BUT NOT MATH. The sum of 1+2+3 = 28. Also that 18 people took math. So 1+2 = 18. But 5 people took both math and lit. So 1 +5 = 18 i.e. no. of people of 1 = 13. So people who took only math but not lit = 13, both math and lit = 5, Now you can subtract sum of 13+5 to find people who took lit but not math

OpenStudy (anonymous):

and that is 18?

OpenStudy (anonymous):

Yep. so the combined sum of 1+2 set = 18. But the number of people who chose something = 28. Hence number of people who took lit, but not math = 28-18

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

10 is the answer. :) (So here, be careful to see "Took math", and "took math ONLY." Both are different. OK?)

OpenStudy (anonymous):

whew! i think i just have to have it completely broken down to me lol thanks!